Use R to plot theoretical ACVF and PACF by hand of the folowing models. (a) MA(3) (b) AR(2) (c) ARMA(2,2)
Answer #1
set.seed(456) ## Simulate MA(3) sample of size 100 ma_sim<-arima.sim(model=list(ma=c(-.2,.1,0.7)),n=100) #Plotting MA2 time setries plot(ma_sim) #Plotting ACF & PACF graph acf(ma_sim) pacf(ma_sim)

## Simulate AR2 sample of size 100 ar_sim<-ts(arima.sim(model=list(ar=c(.9,-.2)),n=100)) #Plotting AR2 time setries plot(ar_sim) #Plotting ACF & PACF graph acf(ar_sim) pacf(ar_sim)

## Simulate ARMA2 sample of size 100 arma_sim<-ts(arima.sim(model=list(ar=c(.9,-.2) , ma=c(-.2,.1)),n=100)) #Plotting ARMA2 time setries plot(arma_sim) #Plotting ACF & PACF graph acf(arma_sim) pacf(arma_sim)